A plot of ln Rate vs. 1/T for a certain reaction gave a straight line with a slope of -8571. What is the activation energy for this reaction? Give your answer in kJ/mole please show work

What is the order of reaction with respect to a certain reactant if doubling the concentration of that reactant causes the reaction rate to double? please explain

What is the order of reaction with respect to a certain reactant if quadrupling the concentration of that reactant causes the rate of reaction to increase by a factor of 16 times? please explain
Answer #1

Slope of curve between ln(rate) versus 1/T is -E/R

So, -E/R = -8571 (R = 8.314 J/K.mol)

E = R×8571 J/mol = 8.314×8571 J/mol = 71259.29 J/mol

= 71.26 KJ/mol (1KJ = 1000J)

Since, on doubling concentration, rate of reaction get doubled thus rate directly depends on concentration of reactant. Hence order is 1.

On quadrupling concentration, rate become 16 times thus rate depend on square of the concentration of reactant. Hence order is 2.

please give feedback!